# Efficient and Effective Adaptation of Multimodal Foundation Models in Sequential Recommendation

### v1

@@ Abstract
Multimodal foundation models (MFMs) have revolutionized sequential recommender systems through advanced representation learning. While Parameter-efficient Fine-tuning (PEFT) is commonly used to adapt these models, studies often prioritize parameter efficiency, neglecting GPU memory and training speed. To address this, we introduced the IISAN framework, significantly enhancing efficiency. However, IISAN was limited to symmetrical MFMs and identical text and image encoders, preventing the use of state-of-the-art Large Language Models. To overcome this, we developed IISAN-Versa, a versatile plug-and-play architecture compatible with both symmetrical and asymmetrical MFMs. IISAN-Versa employs a Decoupled PEFT structure and utilizes both intra- and inter-modal adaptation. It effectively handles asymmetry through a simple yet effective combination of group layer-dropping and dimension transformation alignment. Our research demonstrates that IISAN-Versa effectively adapts large text encoders, and we further identify a scaling effect where larger encoders generally perform better. IISAN-Versa also demonstrates strong versatility in our defined multimodal scenarios, which include raw titles and captions generated from images and videos. Additionally, IISAN-Versa achieved state-of-the-art performance on the Microlens public benchmark. We will release our code and datasets to support future research.
@@ I Introduction
Recent advancements in recommendation algorithms have demonstrated that utilizing powerful large language models (LLMs) and vision encoders like GPT-4o [^1] https://openai.com/index/hello-gpt-4o/, GPT-4 [1], DALL-E [2], Llama [3], and CLIP [4] can obtain a state-of-the-art recommendation performance [5, 6, 7, 8, 9]. These types of models typically combine a recommender model as the user encoder with multimodal foundation models (MFMs) [^2] The definition of multimodal foundation models (MFMs) may vary. In this paper, we focus on separately pre-trained text and image encoders, following the approach in  [10, 4], where the representations of image and text are learned in an end-to-end manner. Other MFM paradigms, such as [11, 12], which use a large language model (LLM) as the backbone and input image encoder features into the LLM, are beyond the scope of this paper. to encode the items. Many studies [5, 6, 7] have shown that fine-tuning the item encoder can lead to optimal performance. However, these approaches are highly inefficient due to the immense computational resources required to fine-tune such large encoders.
@@
A popular paradigm [13] for resolving the efficiency problem includes methods such as Adapter [14], LoRA [15], Bitfit [16], and Compacter [17]. These approaches integrate tunable neural network modules into the backbone multimodal foundation models, collectively referred to as embedded parameter-efficient fine-tuning (EPEFT) [18]. While EPEFT methods have gained popularity for addressing the efficiency of trainable parameters, they still face practical efficiency challenges, such as training time and GPU memory consumption. The introduction of the IISAN [18], with its decoupled structure and caching strategy for multimodal encoders, addresses these practical efficiency issues, outperforming both traditional full fine-tuning and EPEFT methods in terms of both performance and efficiency. We refer to this decoupled, parameter-efficient fine-tuning approach as DPEFT. It significantly reduces GPU memory usage by up to 15 times compared to FFT and 12 times compared to EPEFT (Adapter/LoRA). Additionally, it accelerates training time per epoch by up to 20 times compared to FFT and 16 times compared to EPEFT.
@@
IISAN leverages both intra- and inter-modal information from multimodal foundation models by exploiting the hidden states within the layers of backbone models to achieve fine-grained information merging. Despite these advancements, merging multimodal hidden states from two different encoders presents two limitations: (1) it can only be conducted for symmetrical multi-modal transformer encoders; (2) due to the first limitation, it is hard to explore whether scaling the text encoder in IISAN with recent state-of-the-art LLMs could yield performance improvements. These limitations do not align with the current research trend, where text encoders are often larger and more complex than visual transformers, as shown in Figure 1, which compares the state-of-the-art text and visual transformers. While vision transformers like ViT-e [19], ViT-G [20], and ViT-22B [21] have billions of parameters, they are all closed-source models. Therefore, the most commonly used vision transformers currently have fewer than 1 billion parameters [22]. Normally, larger pre-trained transformers generally offer better performance according to the scaling effect [23]. To maximize the potential of pre-trained models, it is advantageous to use larger models. However, finding a vision transformer that matches the size of a text encoder (LLM) is nearly impossible due to the disparity in model sizes. Therefore, addressing the issue of asymmetrical merging is of paramount importance.
@@
Figure 1: Comparisons of the model sizes of various popular text transformers and visual transformers using a logarithmic scale. The close-sourced is marked as ”Close-sourced” in the figure, the other are all open-sourced models.
@@ II Related Work
Multimodal Foundation Models. Recent work in multimodal learning focuses on leveraging pre-trained models for downstream tasks to enhance performance and reduce pre-training costs [25, 26, 27, 28, 29]. BERT [30] pioneered the pretraining and fine-tuning paradigm in NLP, using Masked Language Modeling and the Next Sentence Prediction to understand context and sentence relationships. Following this, the Vision Transformer (ViT) [22] adapted transformer architecture for image classification, demonstrating competitive performance by treating image patches as tokens. CLIP [4] further bridged vision and language, training a multimodal foundation model to match images with their textual descriptions through contrastive learning, enabling powerful zero-shot learning capabilities [4]. The evolution continued with large-scale models like GPT-4 [1], T5 [31], and multimodal variants such as DALL-E [32] and Flamingo [12]. These models, characterized by their vast scale and diverse training data, perform a wide array of tasks, pushing the boundaries of AI’s multimodal understanding and processing capabilities.
@@
Modality-based Sequential Recommendation (MoRec). The Recommender System community has shown a growing interest in incorporating various modality information into recommendation systems [33, 5, 34, 35, 24]. These systems utilize large-scale multimodal foundation models [30, 36, 4, 34] from NLP and CV [22, 37] to encode text and image data. The sequential encoder remains consistent with traditional architectures like SASRec [38], GRU4Rec [39], and BERT4Rec [40]. Furthermore, models such as IDA-SR [41], UniSRec [42], and VQ-Rec [8] have advanced MoRec by developing item representations using NLP foundation models. For image representation in recommender systems, studies such as [43] and [44] have integrated image features extracted from ResNet-50 [45] into the models.
@@
Furthermore, research [46, 5, 24, 7] has shown that the MoRec framework, applied through end-to-end learning, significantly outperforms previous methods that rely on offline feature extraction. Specifically, [7, 6] emphasized that end-to-end training, which integrates both image and text modalities, significantly surpasses the methods that rely solely on a single modality. Although directly learning the raw content showcases tremendous performance, a significant limitation of these end-to-end studies is their ongoing reliance on fully fine-tuning large-scale multimodal encoders, which often leads to inefficiency.
@@
Parameter-efficient Fine-tuning (PEFT). In the fields of Natural Language Processing (NLP) and Computer Vision (CV), substantial research efforts are focused on investigating Parameter-Efficient Fine-Tuning (PEFT) methods. These methods aim to tackle the challenge posed by the vast number of trainable parameters in large-scale pre-trained models. Foundational works such as Adapter [14], LoRA [15], and BitFit [16] have laid the groundwork in this area. Building on these foundational works, a variety of alternative strategies have emerged, as seen in studies like [17, 47, 48]. Despite the advancements, these approaches primarily utilize Embedded PEFT (EPEFT) and focus largely on parameter efficiency, often overlooking practical efficiency concerns such as training speed and GPU memory usage.
@@
To overcome these limitations, [49] proposed the concept of ”Reduce Memory, Not Parameters,” advocating for a reduction in activations. Additionally, LST [50] offers a memory-efficient strategy aimed at decoupling PEFT modules from the encoder-decoder T5 [31]. A notable shift in recent NLP and CV research is the transition from an emphasis on Embedded PEFT (EPEFT) to Decoupled PEFT (DPEFT), as illustrated by recent works [51, 52, 53]. This shift reflects an evolving understanding that focuses more on practical efficiency instead of trainable parameters.
@@
In modality-based sequential recommender systems (RS), PEFT methods have also shown significant progress, as demonstrated by M6-Rec [54], Tallrec [55], and SSNA [56]. These studies indicate that PEFT approaches can achieve performance levels comparable to traditional fine-tuning methods. However, most such methods still rely on conventional EPEFT, often ignoring practical efficiency aspects. The adoption of DPEFT in recommender systems remains limited, with only exceptions, such as a preprint [56]. Moreover, these methods generally focus on single-modality analysis, missing opportunities to leverage the abundant multimodal information in recommendation scenarios. VIP5 [11] investigates multimodal recommendations using adapters within the P5 framework, though it primarily utilizes adapters for text encoders and leaves image encoders unchanged for feature extraction. This contrasts with our approach, as adapting multiple modality encoders incurs greater costs than only using the text encoder. We initially introduced IISAN [18], an efficient Decoupled PEFT approach designed for symmetrical multimodal foundational models in sequential recommendation. However, it has limitations in effectively handling asymmetrical MFMs, preventing it from leveraging the capabilities of recent state-of-the-art large language models.
@@ III-B Problem Formulation
Given a recommendation dataset $\mathcal{D}=\{\mathcal{U},\mathcal{V}\}$, where $\mathcal{U}$ and $\mathcal{V}$ denote the set of users and the set of items, respectively, we aim to predict the next item interacted with by a user $u$ in a multimodal sequential recommendation task by exploiting their $n$ past behaviors. In multimodal recommendation, each item $v$ contains two modal representations, i.e., text ($v^{\text{text}}$) and corresponding image ($v^{\text{image}}$). These texts and images are fed into two pre-trained multimodal foundation models, such as BERT [30] or Llama [3] for texts and ViT [22] for images, which consist of an embedding layer (word embedding for texts and patch embedding for images) and multiple transformer blocks, respectively. Through these pre-trained multimodal backbones, we obtain textual and visual embeddings from the embedding layers and multiple hidden states ($\{h_{i}^{\text{text}}\}$ and $\{h_{i}^{\text{image}}\}$) from the transformer blocks ($\{TRM_{i}\}$), where $i$ denotes the $i$-th layer of the backbone. Specifically, for IISAN-Versa, we decouple the trainable parameters into three separate towers: the textual-modality side-adapted network for text representation, obtaining the final text embedding $e^{\text{text}}$; the visual-modality side-adapted network for image representation, obtaining the image embedding $e^{\text{image}}$; and an inter-modal side-adapted network for image-text interactive representation, obtaining the inter-modality embedding $e^{\text{inter}}$. A linear-based fusion layer $(FL)$ is added to ensure the consistency of the output dimensions of the item embedding and input dimensions of the sequential encoder following [42, 5] for the final recommendation, as follows:
${e^{item}}=FL([{e^{image}}:{e^{inter}}:{e^{text}}])$ (1)
where [:]delimited-[]:[:][ : ] means the feature concatenation. Then, we input the $\boldsymbol{e^{item}}$ into the sequential encoders and calculate the final predicted score for user $u$ to $i$-th item as $\boldsymbol{\hat{y}_{ui}}$, which is the product of the output sequential encoder and corresponding item embedding.
@@ III-C Intra-model Side Adapted Network
Each side-adapted network (SAN) consists of several SAN blocks (SANBs), which include Upsample and Downsample layers built on a fully connected network. Each SANB, a unit within the SAN, serves as a network block. In this research, we adopted the adapter block as the basic unit of the SAN, termed SANB. Unlike the insertion from Adapter Tuning [14], SANB is implemented in a decoupled manner. For instance, in the textual SAN, a learnable gate mechanism is utilized to fuse the information $h_{i-1}^{B^{intra}}$ from the previous SANB with the current hidden state $h_{i}^{text}$, as demonstrated in the following formula:
$h_{i}^{B^{intra}}=SANB^{intra}\left(\mu_{i}^{text}\ast h_{i-1}^{B^{intra}}+(1-\mu_{i}^{text})\ast h_{i}^{text}\right)$ (2)
where $\mu_{i}^{text}\in[0,1]$. Note that the first SANB will only take the text embeddings and the visual SAN employs similar operations. This Intra-SAN is applicable to both symmetrical and asymmetrical frameworks, the only difference we should make is to enlarge the dimension of input and output dimensions of SANB to fit the larger encoders.
@@ III-D Inter-model Side Adapted Network
When adapting IISAN-Versa to asymmetrical multimodal foundation models, the key difference is in handling the fusion of inter-modality. By performing group layer dropping, we transform the asymmetrical number of hidden states into a symmetrical form, as explained in Section III-E. Following this, we use an intuitive dimension transformation layer (DTL) to align the embedding dimensions of the LLM and the corresponding vision encoders. We then apply a fusion mechanism to merge the hidden states of the two modalities and sum the information from the last SANB using a gating mechanism similar to Intra-SAN, as described by the following formula:
$h_{i}^{B^{inter}}=SANB^{inter}\left(\beta_{i}h_{i}^{image}+(1-\beta_{i})X_{i}+h_{i-1}^{B^{inter}}\right)$ (3)
@@
$X_{i}=\begin{cases}h_{i}^{text},&\text{for IISAN-VS}\\ DTL(h_{i}^{text}),&\text{for IISAN-VA}\end{cases}$ (4)

### v2

@@ Abstract
Multimodal foundation models (MFMs) have revolutionized sequential recommender systems through advanced representation learning. While Parameter-efficient Fine-tuning (PEFT) is commonly used to adapt these models, studies often prioritize parameter efficiency, neglecting GPU memory and training speed. To address this, we introduced the IISAN framework, significantly enhancing efficiency. However, IISAN was limited to symmetrical MFMs and identical text and image encoders, preventing the use of state-of-the-art Large Language Models. To overcome this, we developed IISAN-Versa, a versatile plug-and-play architecture compatible with both symmetrical and asymmetrical MFMs. IISAN-Versa employs a Decoupled PEFT structure and utilizes both intra- and inter-modal adaptation. It effectively handles asymmetry through a simple yet effective combination of group layer-dropping and dimension transformation alignment. Our research demonstrates that IISAN-Versa effectively adapts large text encoders, and we further identify a scaling effect where larger text encoders generally perform better. IISAN-Versa also demonstrates strong versatility in our defined multimodal scenarios, which include raw titles and captions generated from images and videos. Additionally, IISAN-Versa achieved state-of-the-art performance on the MicroLens public benchmark. We release our code at https://github.com/GAIR-Lab/IISAN.
@@ I Introduction
Recent advancements in recommendation algorithms have demonstrated that utilizing powerful large language models (LLMs) and vision encoders like GPT, Llama [^1] https://www.llama.com/, https://openai.com/index/hello-gpt-4o/, and CLIP [1] can obtain a state-of-the-art recommendation performance [2, 3, 4, 5, 6]. These types of models typically combine a recommender model as the user encoder with multimodal foundation models (MFMs) [^2] The definition of multimodal foundation models (MFMs) may vary. In this paper, we focus on separately pre-trained text and image encoders, following the approach in  [7, 1], where the representations of image and text are learned in an end-to-end manner. Other MFM paradigms, such as [8, 9], which use a large language model (LLM) as the backbone and input image encoder features into the LLM, are beyond the scope of this paper. to encode the items. Many studies [2, 3, 4] have shown that fine-tuning the item encoder can lead to optimal performance. However, these approaches are highly inefficient due to the immense computational resources required to fine-tune such large encoders.
@@
A popular paradigm [10] for resolving the efficiency problem includes methods such as Adapter [11] and LoRA [12], BitFit [13]. These approaches integrate tunable neural network modules into the backbone multimodal foundation models, collectively referred to as embedded parameter-efficient fine-tuning (EPEFT) [14]. While EPEFT methods have gained popularity for addressing the efficiency of trainable parameters, they still face practical efficiency challenges, such as training time and GPU memory consumption. The introduction of the IISAN [14], with its decoupled structure and caching strategy for multimodal encoders, addresses these practical efficiency issues, outperforming both traditional full fine-tuning and EPEFT methods in terms of both performance and efficiency. We refer to this decoupled, parameter-efficient fine-tuning approach as DPEFT. It significantly reduces GPU memory usage by up to 15 times compared to FFT and 12 times compared to EPEFT (Adapter/LoRA). Additionally, it accelerates training time per epoch by up to 20 times compared to FFT and 16 times compared to EPEFT.
@@
IISAN leverages both intra- and inter-modal information from multimodal foundation models by exploiting the hidden states within the layers of backbone models to achieve fine-grained information merging. Despite these advancements, merging multimodal hidden states from two different encoders presents two limitations: (1) it can only be conducted for symmetrical multimodal transformer encoders; (2) due to the first limitation, it is hard to explore whether scaling the text encoder in IISAN with recent state-of-the-art LLMs could yield performance improvements. These limitations do not align with the current research trend, where text encoders are often larger and more complex than visual transformers, as shown in Figure 1, which compares the state-of-the-art text and visual transformers. While vision transformers like ViT-22B [15] have billions of parameters, they are all closed-source models. Therefore, the most commonly used vision transformers currently have fewer than 1 billion parameters [16]. Normally, larger pre-trained transformers generally offer better performance according to the scaling effect [17]. To maximize the potential of pre-trained models, it is advantageous to use larger models. However, finding a vision transformer that matches the size of a text encoder (LLM) is nearly impossible due to the disparity in model sizes. Therefore, addressing the issue of asymmetrical merging is of paramount importance.
@@
Figure 1: Comparisons of model sizes for various popular text transformers (T-TRM) and visual transformers (V-TRM) on a logarithmic scale. Closed-source models are indicated by “(Closed)” in the figure; all other models are open-source.
@@ II Related Work
Multimodal Foundation Models. Recent work in multimodal learning focuses on leveraging pre-trained models for downstream tasks to enhance performance and reduce pre-training costs [19]. BERT [20] pioneered the pretraining and fine-tuning paradigm in NLP, using Masked Language Modeling and the Next Sentence Prediction to understand context and sentence relationships. Following this, the Vision Transformer (ViT) [16] adapted transformer architecture for image classification, demonstrating competitive performance by treating image patches as tokens. CLIP [1] further bridged vision and language, training a multimodal foundation model to match images with their textual descriptions through contrastive learning, enabling powerful zero-shot learning capabilities [1]. The evolution continued with large-scale models like GPT-4, and multimodal variants such as Flamingo [9]. These models, characterized by their vast scale and diverse training data, perform a wide array of tasks, pushing the boundaries of AI’s multimodal understanding and processing capabilities.
@@
Modality-based Sequential Recommendation (MoRec). The Recommender System community has shown a growing interest in incorporating various modality information into recommendation systems [21, 2, 22, 18, 23]. These systems utilize large-scale multimodal foundation models [20, 1, 24] from NLP and CV [16, 25] to encode text and image data. The sequential encoder remains consistent with traditional architectures like SASRec [26], GRU4Rec [27], and BERT4Rec [28]. Furthermore, models such as IDA-SR [29], UniSRec [30], and VQ-Rec [5] have advanced MoRec by developing item representations using NLP foundation models. For image representation in recommender systems, studies such as [31] and [32] have integrated image features extracted from ResNet-50 [33] into the models.
@@
Furthermore, research [2, 18, 4] has shown that the MoRec framework, applied through end-to-end learning, significantly outperforms previous methods that rely on offline feature extraction. Specifically, [4, 3] emphasized that end-to-end training, which integrates both image and text modalities, significantly surpasses the methods that rely solely on a single modality. Although directly learning the raw content showcases tremendous performance, a significant limitation of these end-to-end studies is their ongoing reliance on fully fine-tuning large-scale multimodal encoders, which often leads to inefficiency.
@@
Parameter-efficient Fine-tuning (PEFT). In the fields of Natural Language Processing (NLP) and Computer Vision (CV), substantial research efforts are focused on investigating Parameter-Efficient Fine-Tuning (PEFT) methods. These methods aim to tackle the challenge posed by the vast number of trainable parameters in large-scale pre-trained models. Foundational works such as Adapter [11], LoRA [12], and BitFit [13] have laid the groundwork in this area. Building on these foundational works, a variety of alternative strategies have emerged, as seen in studies like [34]. Despite the advancements, these approaches primarily utilize Embedded PEFT (EPEFT) and focus largely on parameter efficiency, often overlooking practical efficiency concerns such as training speed and GPU memory usage.
@@
To overcome these limitations, [35] proposed the concept of Reduce Memory, Not Parameters,” advocating for a reduction in activations. Additionally, LST [36] offers a memory-efficient strategy aimed at decoupling PEFT modules. A notable shift in recent NLP and CV research is the transition from an emphasis on Embedded PEFT (EPEFT) to Decoupled PEFT (DPEFT), as illustrated by recent works [37, 38, 39]. This shift reflects an evolving understanding that focuses more on practical efficiency instead of trainable parameters.
@@
In modality-based sequential recommender systems (RS), PEFT methods have also shown significant progress, as demonstrated by M6-Rec [40], and Tallrec [41]. These studies indicate that PEFT approaches can achieve performance levels comparable to traditional fine-tuning methods. However, most such methods still rely on conventional EPEFT, often ignoring practical efficiency aspects. The adoption of DPEFT in recommender systems remains limited. Moreover, these methods generally focus on single-modality analysis, missing opportunities to leverage the abundant multimodal information in recommendation scenarios. VIP5 [8] investigates multimodal recommendations using adapters within the P5 framework, though it primarily utilizes adapters for text encoders and leaves image encoders unchanged for feature extraction. This contrasts with our approach, as adapting multiple modality encoders incurs greater costs than only using the text encoder. We initially introduced IISAN [14], an efficient Decoupled PEFT approach designed for symmetrical multimodal foundational models in sequential recommendation. However, it has limitations in effectively handling asymmetrical MFMs, preventing it from leveraging the capabilities of recent state-of-the-art large language models.
@@ III-B Problem Formulation
Given a recommendation dataset $\mathcal{D}=\{\mathcal{U},\mathcal{V}\}$, where $\mathcal{U}$ and $\mathcal{V}$ denote the set of users and the set of items, respectively, we aim to predict the next item interacted with by a user $u$ in a multimodal sequential recommendation task by exploiting their $n$ past behaviors. In multimodal recommendation, each item $v$ contains two modal representations, i.e., text ($v^{\text{t}}$) and corresponding image ($v^{\text{v}}$). These texts and images are fed into two pre-trained multimodal foundation models, such as BERT [20] or Llama [42] for texts and ViT [16] for images, which consist of an embedding layer (word embedding for texts and patch embedding for images) and multiple transformer blocks, respectively. Through these pre-trained multimodal backbones, we obtain textual and visual embeddings from the embedding layers and multiple hidden states ($\{h_{i}^{\text{t}}\}$ and $\{h_{i}^{\text{v}}\}$) from the transformer blocks ($\{TRM_{i}\}$), where $i$ denotes the $i$-th layer of the backbone. Specifically, for IISAN-Versa, we decouple the trainable parameters into three separate towers: the textual-modality side-adapted network for text representation, obtaining the final text embedding $e^{\text{t}}$; the visual-modality side-adapted network for image representation, obtaining the image embedding $e^{\text{v}}$; and an inter-modal side-adapted network for image-text interactive representation, obtaining the inter-modality embedding $e^{\text{inter}}$. A linear-based fusion layer $(FL)$ is added to ensure the consistency of the output dimensions of the item embedding and input dimensions of the sequential encoder following [30, 2] for the final recommendation, as follows:
${e^{item}}=FL([{e^{v}}:{e^{inter}}:{e^{t}}])$ (1)
where $[:]$ means the feature concatenation. Then, we input the $\boldsymbol{e^{item}}$ into the sequential encoders and calculate the final predicted score for user $u$ to $i$-th item as $\boldsymbol{\hat{y}_{ui}}$, which is the product of the output sequential encoder and corresponding item embedding.
@@ III-C Intra-model Side Adapted Network
Each side-adapted network (SAN) consists of several SAN blocks (SANBs), which include Upsample and Downsample layers built on a fully connected network. Each SANB, a unit within the SAN, serves as a network block. In this research, we adopted the adapter block as the basic unit of the SAN, termed SANB. Unlike the insertion from Adapter Tuning [11], SANB is implemented in a decoupled manner. For instance, in the textual SAN, a learnable gate mechanism is utilized to fuse the information $h_{i-1}^{B^{intra}}$ from the previous SANB with the current hidden state $h_{i}^{t}$, as demonstrated in the following formula:
$h_{i}^{B^{intra}}=SANB^{intra}\left(\mu_{i}^{t}\ast h_{i-1}^{B^{intra}}+(1-\mu_{i}^{t})\ast h_{i}^{t}\right)$ (2)
where $\mu_{i}^{t}\in[0,1]$. Note that the first SANB will only take the text embeddings and the visual SAN employs similar operations. This Intra-SAN is applicable to both symmetrical and asymmetrical frameworks, the only difference we should make is to enlarge the dimension of input and output dimensions of SANB to fit the larger encoders.
@@ III-D Inter-model Side Adapted Network
When adapting IISAN-Versa to asymmetrical multimodal foundation models, the key difference is in handling the fusion of inter-modality. By performing group layer dropping, we transform the asymmetrical number of hidden states into a symmetrical form, as explained in Section III-E. Following prior work [10, 2], we adopt a Dimension Transformation Layer (DTL) to align the high-dimensional embeddings from the LLM with the lower-dimensional outputs of vision encoders. The DTL is computed as:
$\mathbf{z}=\mathbf{W}\mathbf{x}+\mathbf{b},\quad\mathbf{W}\in\mathbf{R}^{d_{\text{out}}\times d_{\text{in}}},\ \mathbf{b}\in\mathbf{R}^{d_{\text{out}}}$
where $\mathbf{x}\in\mathbf{R}^{d_{\text{in}}}$ is the input feature, $\mathbf{z}\in\mathbf{R}^{d_{\text{out}}}$ is the transformed output, $d_{\text{in}}$ and $d_{\text{out}}$ denote the input and output dimensions respectively, and $d_{\text{in}}>d_{\text{out}}$.
We then apply a fusion mechanism to merge the hidden states of the two modalities and sum the information from the last SANB using a gating mechanism similar to Intra-SAN, as described by the following formula:
$h_{i}^{B^{inter}}=SANB^{inter}\left(\beta_{i}h_{i}^{v}+(1-\beta_{i})X_{i}+h_{i-1}^{B^{inter}}\right)$ (3)
@@
$X_{i}=\begin{cases}h_{i}^{t},&\text{for IISAN-VS}\\ DTL(h_{i}^{t}),&\text{for IISAN-VA}\end{cases}$ (4)
Algorithm 1 Group LayerDrop for layer selection
1:function SelectLayers($L,L^{v}$)
2:  $step\leftarrow\lceil L/L^{v}\rceil$
3:  $S\leftarrow[\,]$
4:  for $i=L^{v}-1$ to 0 do
5:   append $L-step\times i$ to $S$
6:  end for
7:  return $S$
8:end function
TABLE I: Selected layers $S$ for various total depths $L$ (with $L^{v}=6$).
Number of Layers Selected layers
12 $\{2,4,6,8,10,12\}$
24 $\{4,8,12,16,20,24\}$
28 $\{3,8,13,18,23,28\}$
32 $\{2,8,14,20,26,32\}$
48 $\{3,12,21,30,39,48\}$
80 $\{5,20,35,50,65,80\}$